When potassium chlorate is heated decomposes to produce potassium chloride and oxygen gas right advance equation for this reaction and calculate the mass of potassium chlorate needed to produce 5.00x10 raised to the second power milliliters of oxygen gas at 1.10 ATM and 39.0°C you think figs to answer this question
1
Expert's answer
2019-05-06T05:05:31-0400

According to combined gas law:


"PV/T=P0V0/T0 \implies V0=PVT0/P0T"

According to IUPAC:

P0=1ATM

T0=273.15K

P=1.1ATM

T=39oC=39+273.15=312.15K

V(O2)=5.00*102ml=0.5L

V0(O2)= 0.48L

Accordong to Avogadro's hypothesis, which states: Equal volumes of gases at the same temperature and pressure contain equal numbers of molecules. From this, we derive the molar volume of a gas (volume/moles of gas):

Vm=22.414L/mol


"Vm=V0/n\implies n=V0/Vm"


n=0.48/22.414=0.02mol

With a respect to molar ratios:

n(KClO3)=0.02*2/3=0.01mol

m(KClO3)=M(KClO3)*n(KClO3)

M(KClO3)=39+35.5+16*3=122.5g/mol

m(KClO3)=122.5*0.01=1.22g

Answer: mass of potassium chlorate is 1.22g